# thanks: Corresponding author

### v1

@@ II Theory and Computational Details
While our numerical evidence clearly demonstrates that the CIPSI computations characterized by $N_{\text{det(in)}}=10,000,000$ result in the $P$ spaces which are unnecessarily large for accurately approximating the CCSDT energetics using the CIPSI-driven CC($P$;$Q$) approach, we include them in our analysis since they also allowed us to extrapolate the near-full-CI $E_{\text{var}}+\Delta E_{r}^{(2)}$ potentials shown in Figure 1 by following the procedure described in Refs. Garniron et al., 2019; Loos et al., 2020; Monino et al., 2022 (see, also, Ref. Gururangan et al., 2021). In this procedure, the $E_{\text{var},k}+\Delta E_{r,k}^{(2)}$ energies extracted from the last four to six Hamiltonian diagonalizations of the CIPSI sequence leading to the final $|\Psi^{(\text{CIPSI})}\rangle$ state are plotted against the corresponding $\Delta E_{r,k}^{(2)}$ corrections and the resulting data, fit to a line, are extrapolated to the $\Delta E_{r}^{(2)}=0$ limit. In the case of the CIPSI calculations performed in this work, to extrapolate the reasonably smooth PECs for the lowest singlet and triplet states of cyclobutadiene out of our largest CIPSI runs corresponding to $N_{\text{det(in)}}=10,000,000$, shown in Figure 1, we had to use the last six $E_{\text{var},k}+\Delta E_{r,k}^{(2)}$ and $\Delta E_{r,k}^{(2)}$ values of each of these runs. Using fewer than six values resulted in unphysical bumps in the extrapolated potentials. In addition to allowing us to comment on the quality of the CCSDT energetics in the Introduction, the extrapolated CIPSI PECs shown in Figure 1 serve in this study as reference potentials for assessing the accuracy of the $E_{\text{var}}$, $E_{\text{var}}+\Delta E^{(2)}$, and $E_{\text{var}}+\Delta E_{r}^{(2)}$ energies obtained in the CIPSI calculations using various values of $N_{\text{det(in)}}$ from a 1–10,000,000 range.
In generating the results of the CIPSI-driven CC($P$;$Q$) computations for the lowest singlet and triplet states of cyclobutadiene along its automerization coordinate at different values of the wave function terminating parameter $N_{\text{det(in)}}$, discussed in the next section, we adopted the same philosophy as that used in the case of the DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$, extrapolated $E_{\text{var}}+\Delta E_{r}^{(2)}$, and parent CCSDT potentials shown in Figure 1. Thus, we used Eq. (2), in which the geometries of the R and TS structures on the singlet potential obtained with MR-AQCC were taken from Ref. Eckert-Maksić et al., 2006, to set up a one-dimensional, $D_{\rm 2h}$-symmetric, automerization pathway parameterized by dimensionless variable $\lambda\in[0,1]$. In analogy to the PECs shown in Figure 1, we carried out our CIPSI-based CC($P$;$Q$) and the underlying CC($P$) calculations for the lowest ${}^{1}A_{g}(D_{2\text{h}})$ and ${}^{3}B_{1g}(D_{2\text{h}})$ states of cyclobutadiene and determined the corresponding $E_{\text{var}}$, $E_{\text{var}}+\Delta E^{(2)}$, and $E_{\text{var}}+\Delta E_{\text{r}}^{(2)}$ CIPSI energies for $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1, reflecting the resulting PEC segments about $\lambda=1$ to obtain the potentials that connect the rectangular reactant and product minima via the square TS. As in the case of the full CCSDT, perturbatively corrected and extrapolated CIPSI, and DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$ calculations reported in Figure 1, all of our CIPSI-driven CC($P$) and CC($P$;$Q$) computations were performed using the cc-pVDZ basis set and the four core molecular orbitals correlating with the 1s shells of the carbon atoms were frozen in the post-RHF/ROHF steps. Consistent with the overall symmetry of the automerization pathway examined in this work, the $D_{\rm 2h}$ point group was adopted throughout. In particular, the $P$ and $Q$ spaces used in our CC($P$;$Q$) calculations for the lowest-energy singlet PEC consisted of the $S_{z}=0$ determinants of the $A_{g}(D_{\rm 2h})$ symmetry. In the case of the lowest-energy triplet potential, they consisted of the $S_{z}=1$ $B_{1g}(D_{\rm 2h})$ determinants.
@@ III Results and Discussion
As explained in the Introduction, the primary objective of this study is to examine efficiency of the CIPSI-driven CC($P$;$Q$) methodology in converging the full CCSDT data for the lowest singlet and triplet potentials of cyclobutadiene along its automerization coordinate and the gap between them. We are especially interested in investigating how effective the CIPSI-driven CC($P$;$Q$) approach is in balancing the substantial nondynamical many-electron correlation effects characterizing the lowest ${}^{1}A_{g}(D_{\rm 2h})$ state in the neighborhood of the automerization barrier region, where $T_{3}$ clusters become large, nonperturbative, and strongly coupled to their lower-rank $T_{1}$ and $T_{2}$ counterparts, with the predominantly dynamical correlations characterizing the lowest ${}^{3}B_{1g}(D_{\rm 2h})$ state. As pointed out in the Introduction, using comparisons with the DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$ and perturbatively corrected and extrapolated CIPSI results, the CCSDT approach, despite its intrinsically single-reference character, captures essentially all relevant many-electron correlation effects needed to accurately describe the lowest singlet and triplet potentials of cyclobutadiene and the separation between them. The question is if the CIPSI-driven CC($P$;$Q$) computations using compact wave functions $|\Psi^{(\text{CIPSI})}\rangle$, resulting from the relatively inexpensive CIPSI diagonalization sequences characterized by the $N_{\text{det(out)}}$ values that are much smaller than the numbers of all $T_{3}$ amplitudes, and tiny fractions of the triply excited determinants in the underlying $P$ spaces are capable of accomplishing the same. We also examine how effective the noniterative $\delta(\mbox{$P$;$Q$})$ corrections are in accelerating convergence of the CC($P$) energetics toward their CCSDT parents and how the rate of convergence of the CIPSI-driven CC($P$;$Q$) calculations toward CCSDT with $N_{\text{det(in)}}$ compares to the analogous convergence of the perturbatively corrected CIPSI energies toward their extrapolated $E_{\text{var}}+\Delta E_{\text{r}}^{(2)}$ values. The results of our CIPSI-driven CC($P$) and CC($P$;$Q$) calculations for the lowest ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ potentials and $\Delta E_{\text{S--T}}$ gaps of cyclobutadiene, along with the associated $E_{\text{var}}$, $E_{\text{var}}+\Delta E^{(2)}$, and $E_{\text{var}}+\Delta E^{(2)}_{r}$ data, can be found in Tables 13. The CC($P$) and CC($P$;$Q$) potentials and $\Delta E_{\text{S--T}}$ values are also shown in Figure 2.
@@
We begin by commenting on the calculations using $N_{\text{det(in)}}=1$, which help us appreciate the need for including the leading triply excited determinants in the $P$ spaces employed in the CC($P$) and CC($P$;$Q$) computations, especially when $T_{3}$ effects and the coupling between $T_{1}$ and $T_{2}$ clusters and their higher-rank $T_{3}$ counterpart become significant. As explained in Section II, when $N_{\text{det(in)}}=1$, the CIPSI-driven CC($P$) and CC($P$;$Q$) approaches become equivalent to CCSD and CR-CC(2,3), respectively, i.e., one solves the CCSD equations for $T_{1}$ and $T_{2}$ clusters, as if they were decoupled from their higher-rank $T_{3}$ counterpart, and corrects the resulting CCSD energies for the effects of connected triples using the CR-CC(2,3) method. Upon examining the $N_{\text{det(in)}}=1$ CC($P$) values in Table 1, we observe that the $T_{3}$ correlation effects characterizing the lowest-energy ${}^{1}A_{g}(D_{2\text{h}})$ potential, estimated by forming differences between the CCSDT and CCSD energies, are not only large, but also dramatically changing along the automerization coordinate, from $-26.827$ millihartree for the $\lambda=0$ R species to $-47.979$ millihartree when the square TS structure corresponding to $\lambda=1$ is considered. As indicated by the $N_{\text{det(in)}}=1$ CC($P$;$Q$) results shown in Table 1, the incorporation of $T_{3}$ correlations using the noniterative CR-CC(2,3) corrections to the CCSD energies helps, reducing the 26.827, 27.964, 29.667, 32.473, 37.662, and 47.979 millihartree errors relative to CCSDT obtained at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1 with CCSD to 0.848, 1.253, 2.021, 3.582, 7.008, and 14.636 millihartree, respectively, but substantial discrepancies between the CR-CC(2,3) and CCSDT data, especially in the neighborhood of the barrrier region, where they are as large as 7–15 millihartree when $\lambda\in[0.8,1]$, remain. As a result, the quality of the ${}^{1}A_{g}(D_{2\text{h}})$ potential obtained in the CR-CC(2,3) calculations, which is characterized by a large, 13.788 millihartree, nonparallelity error (NPE) relative to its CCSDT parent and which is shown in Figure 2 (b), is poor. Failure of CR-CC(2,3) and, as demonstrated, for example, in Refs. Shen and Piecuch, 2012; Lyakh et al., 2011, of other noniterative triples corrections to CCSD, including CCSD(T)Shen and Piecuch (2012); Lyakh et al. (2011) and CCSD(2)T,Shen and Piecuch (2012) to accurately describe the lowest ${}^{1}A_{g}(D_{2\text{h}})$ state of cyclobutadiene in the automerization barrier region is, in significant part, a consequence of the inability of all such methods to capture the coupling of the lower-rank $T_{1}$ and $T_{2}$ clusters with $T_{3}$, which in the vicinity of the TS geometry, where $T_{3}$ effects are large and nonperturbative, is big enough to substantially alter $T_{1}$ and $T_{2}$ amplitudes compared to their CCSD values. This means that to improve the quality of the ${}^{1}A_{g}(D_{2\text{h}})$ potential obtained with CR-CC(2,3), one should relax $T_{1}$ and $T_{2}$ clusters, adjusting them to the presence of $T_{3}$ correlations, prior to determining noniterative triples corrections. The CIPSI-driven CC($P$;$Q$) methodology allows us to do it in a computationally efficient manner, avoiding expensive CCSDT iterations, by incorporating the leading triply excited determinants identified with the help of the CIPSI runs using sufficiently large $N_{\text{det(in)}}>1$ values into the underlying $P$ spaces and correcting the resulting CC($P$) energies for the remaining $T_{3}$ effects using the $\delta(\mbox{$P$;$Q$})$ corrections.
@@
The situation with the lowest-energy ${}^{3}B_{1g}(D_{\rm 2h})$ potential is different. In this case, as shown in Table 2, the $T_{3}$ correlation effects, estimated, after subtracting the CCSD energies from their CCSDT counterparts, at about $(-25)-(-24)$ millihartree, barely vary with the automerization coordinate $\lambda$ and are accurately described by the CR-CC(2,3) approach, which reproduces the parent CCSDT energetics to within 60 microhartree across the entire ${}^{3}B_{1g}(D_{\rm 2h})$ PEC. Because of this very different behavior of CR-CC(2,3) compared to the lowest-energy singlet potential, which can be seen by comparing the $N_{\text{det(in)}}=1$ CC($P$;$Q$) PECs in the (b) and (d) panels of Figure 2, one ends up with a highly unbalanced description of the lowest singlet and triplet states of cyclobutadiene by the CR-CC(2,3) method, especially in vicinity of the barrier on the ${}^{1}A_{g}(D_{2\text{h}})$ potential. This is reflected in the errors characterizing the singlet–triplet gap values obtained with CR-CC(2,3), relative to CCSDT, which grow from $0.553$ kcal/mol at $\lambda=0$ to $9.222$ kcal/mol when the $\lambda=1$ square TS structure is considered [see the $N_{\text{det(in)}}=1$ CC($P$;$Q$) values of $\Delta E_{\text{S--T}}$ in Table 3; cf., also, Figure 2 (f)]. Once again, the main problem resides in the neglect of the coupling between $T_{1}$ and $T_{2}$ clusters and their higher-rank $T_{3}$ counterpart in the CR-CC(2,3) approach, which results in a poor description of the ${}^{1}A_{g}(D_{2\text{h}})$ potential in the vicinity of the TS geometry that propagates into the similarly poor $\Delta E_{\text{S--T}}$ gap values. To bring the results closer to those obtained with CCSDT, the input variable $N_{\text{det(in)}}$ that controls the CIPSI runs preceding the CC($P$) and CC($P$;$Q$) steps must be increased. The results of the CIPSI-driven CC($P$) and CC($P$;$Q$) computations for the lowest ${}^{1}A_{g}(D_{2\text{h}})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ potentials of cyclobutadiene and the gap between them using representative $N_{\text{det(in)}}>1$ values are discussed next.
@@
As shown in Table 1 and Figure 2 (b), the CC($P$;$Q$) computations for the lowest ${}^{1}A_{g}(D_{2\text{h}})$ state, using CIPSI Hamiltonian diagonalizations to identify the leading triply excited determinants for inclusion in the underlying $P$ spaces, display fast convergence toward CCSDT with $N_{\text{det(in)}}$, independent of the value of $\lambda$. With as little as 101,361101361101,361101 , 361196,965196965196,965196 , 965 $S_{z}=0$ determinants of the $A_{g}(D_{\rm 2h})$ symmetry in the terminal $|\Psi^{(\text{CIPSI})}\rangle$ wave functions generated by the inexpensive CIPSI runs using $N_{\text{det(in)}}=100,000$, which capture tiny fractions, on the order of 0.1–0.2%, of the $14,483,876$ $A_{g}(D_{\rm 2h})$-symmetric $S_{z}=0$ triples, the CC($P$;$Q$) method reduces the 0.848, 1.253, 2.021, 3.582, 7.008, and 14.636 millihartree errors relative to CCSDT obtained at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1 with CR-CC(2,3) to 0.431, 0.552, 0.776, 1.235, 0.628, and 3.539 millihartree, respectively. With the relatively small additional effort corresponding to $N_{\text{det(in)}}=250,000$, which results in 394,080394080394,080394 , 080449,753449753449,753449 , 753 $S_{z}=0$ determinants of the $A_{g}(D_{\rm 2h})$ symmetry in the final CIPSI diagonalization spaces and only 0.5–0.6% of all triples in the underlying $P$ spaces, the differences between the CC($P$;$Q$) and CCSDT energies of the lowest singlet state of cyclobutadiene at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1 decrease to 0.278, 0.272, 0.271, 0.300, 0.336, and 0.458 millihartree, respectively. Clearly, these are massive error reductions compared to the CR-CC(2,3) computations, especially in the barrier region, which highlight the effectiveness of our CIPSI-driven CC($P$;$Q$) strategy and the importance of relaxing $T_{1}$ and $T_{2}$ amplitudes in the presence of the leading $T_{3}$ contributions compared to their CCSD values prior to determining the noniterative corrections for the remaining $T_{3}$ effects. Similar comments apply to the improvements in the troublesome 13.788 millihartree NPE relative to CCSDT characterizing the ${}^{1}A_{g}(D_{2\text{h}})$ potential obtained in the CR-CC(2,3) calculations offered by the CIPSI-driven CC($P$;$Q$) runs. When the CC($P$;$Q$) approach using $N_{\text{det(in)}}=100,000$ is employed, the NPE relative to CCSDT characterizing the resulting ${}^{1}A_{g}(D_{2\text{h}})$ potential becomes 3.108 millihartree, which is a reduction of the CR-CC(2,3) NPE value by a factor of 4.4. The $N_{\text{det(in)}}=250,000$ CC($P$;$Q$) computations, which extract the lists of triples from the relatively inexpensive Hamiltonian diagonalizations in spaces that are 32–37 times smaller than the number of $T_{3}$ amplitudes used by CCSDT, reduce the 13.788 millihartree NPE characterizing the lowest-energy ${}^{1}A_{g}(D_{2\text{h}})$ potential obtained with CR-CC(2,3), relative to its CCSDT counterpart, to an impressively small value of 0.187 millihartree. This is a 74-fold reduction in NPE compared to CR-CC(2,3).

### v2

@@ II Theory and Computational Details
While the numerical evidence discussed in the next section clearly demonstrates that the CIPSI computations characterized by $N_{\text{det(in)}}=10,000,000$ result in the $P$ spaces which are unnecessarily large for accurately approximating the CCSDT energetics using the CIPSI-driven CC($P$;$Q$) approach when the cc-pVDZ and cc-pVTZ basis sets are employed, we include them in our analysis since they also allowed us to extrapolate the near-full-CI $E_{\text{var}}+\Delta E_{r}^{(2)}$ potentials, such as those presented in Figure 1 for a cc-pVDZ basis, by following the procedure described in Refs. Garniron et al., 2019; Loos et al., 2020; Monino et al., 2022 (see, also, Ref. Gururangan et al., 2021). In this procedure, the $E_{\text{var},k}+\Delta E_{r,k}^{(2)}$ energies extracted from the last four to six Hamiltonian diagonalizations of the CIPSI sequence leading to the final $|\Psi^{(\text{CIPSI})}\rangle$ state are plotted against the corresponding $\Delta E_{r,k}^{(2)}$ corrections and the resulting data, fit to a line, are extrapolated to the $\Delta E_{r}^{(2)}=0$ limit. In the case of the CIPSI calculations employing the cc-pVDZ basis set, to extrapolate the reasonably smooth PECs for the lowest singlet and triplet states of cyclobutadiene out of our largest CIPSI runs corresponding to $N_{\text{det(in)}}=10,000,000$, shown in Figure 1, we used the last six $E_{\text{var},k}+\Delta E_{r,k}^{(2)}$ and $\Delta E_{r,k}^{(2)}$ values of each of these runs, since using fewer than six values resulted in unphysical bumps in the extrapolated potentials. The determination of the extrapolated $E_{\text{var}}+\Delta E_{r}^{(2)}$ energies of the lowest $^{1}A_{g}(D_{2\text{h}})$ and $^{3}B_{1g}(D_{2\text{h}})$ states of cyclobutadiene at the R and TS geometries using the cc-pVTZ basis had to be handled differently. In this case, we relied on the last four $E_{\text{var},k}+\Delta E_{r,k}^{(2)}$ and $\Delta E_{r,k}^{(2)}$ values obtained in the $N_{\text{det(in)}}=10,000,000$ CIPSI runs. Using more values than four led to problems with producing a sensible result for the extrapolated $E_{\text{var}}+\Delta E_{r}^{(2)}$ energy of the lowest singlet state at the TS geometry. In addition to allowing us to comment on the quality of the CCSDT energetics in the Introduction, the extrapolated $E_{\text{var}}+\Delta E_{r}^{(2)}$ energies serve in this study as the reference data for assessing the accuracy of the $E_{\text{var}}$, $E_{\text{var}}+\Delta E^{(2)}$, and $E_{\text{var}}+\Delta E_{r}^{(2)}$ values obtained in the CIPSI calculations using various choices of $N_{\text{det(in)}}$ from a 1–10,000,000 range.
@@ III Results and Discussion
[…]
We begin by analyzing our CIPSI-driven CC($P$) and CC($P$;$Q$) computations employing the cc-pVDZ basis set, used in most of the calculations reported in this work. The results of our CIPSI-driven CC(PP)/$c$-pVDZ and CC($P$;QQ)/$c$-pVDZ calculations for the lowest-energy $^{1}A_{g}(D_{\rm 2h})$ and $^{3}B_{1g}(D_{\rm 2h})$ potentials of cyclobutadiene along its automerization coordinate, the separation between them, and the associated $E_{\text{var}}$, $E_{\text{var}}+\Delta E^{(2)}$, and $E_{\text{var}}+\Delta E^{(2)}_{r}$ data can be found in Tables 13. The convergence of the CC(PP)/$c$-pVDZ and CC($P$;QQ)/$c$-pVDZ energies of the lowest singlet and triplet states of cyclobutadiene and the gaps between them, determined at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1, toward their CCSDT/cc-pVDZ parents as functions of the parameters $N_{\text{det(in)}}$ and $N_{\text{det(out)}}$ that control $N_{\text{det(in)}}$] and define $N_{\text{det(out)}}$] the sizes of the terminal wave functions $|\Psi^{(\text{CIPSI})}\rangle$ produced by the underlying CIPSI runs is also visualized in Figures 24 and, in a summary form, in Figure 5.
@@
Let us first comment on the calculations using $N_{\text{det(in)}}=1$, which help us appreciate the need for including the leading triply excited determinants in the $P$ spaces employed in the CC($P$) and CC($P$;$Q$) computations, especially when $T_{3}$ effects and the coupling between $T_{1}$ and $T_{2}$ clusters and their higher-rank $T_{3}$ counterpart become significant. As explained in Section II, when $N_{\text{det(in)}}=1$, the CIPSI-driven CC($P$) and CC($P$;$Q$) approaches become equivalent to CCSD and CR-CC(2,3), respectively, i.e., one solves the CCSD equations for $T_{1}$ and $T_{2}$ clusters, as if they were decoupled from their higher-rank $T_{3}$ counterpart, and corrects the resulting CCSD energies for the effects of connected triples using the CR-CC(2,3) method. Upon examining the $N_{\text{det(in)}}=1$ CC($P$) values in Table 1 (cf., also, Figure 2), we observe that the $T_{3}$ correlation effects characterizing the lowest-energy $^{1}A_{g}(D_{2\text{h}})$ potential, estimated by forming differences between the CCSDT and CCSD energies, are not only large, but also dramatically changing along the automerization coordinate, from 26.827-26.827 millihartree for the $\lambda=0$ R species to 47.979-47.979 millihartree for the $\lambda=1$ TS structure when the cc-pVDZ basis set is employed. As indicated by the $N_{\text{det(in)}}=1$ CC($P$;$Q$) results shown in Table 1 and Figure 2, the incorporation of $T_{3}$ correlations using the noniterative CR-CC(2,3) corrections to the CCSD energies helps, reducing the 26.827, 27.964, 29.667, 32.473, 37.662, and 47.979 millihartree errors relative to CCSDT obtained at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1 with CCSD to 0.848, 1.253, 2.021, 3.582, 7.008, and 14.636 millihartree, respectively, but substantial discrepancies between the CR-CC(2,3) and CCSDT data, especially in the neighborhood of the barrier region, where they are as large as 7–15 millihartree when $\lambda\in[0.8,1]$, remain. As a result, the quality of the $^{1}A_{g}(D_{2\text{h}})$ potential obtained in the CR-CC(2,3) calculations, which is characterized by a large, 13.788 millihartree, nonparallelity error (NPE) relative to its CCSDT parent and which is shown in Figure 5 (b), is poor. Failure of CR-CC(2,3) and, as demonstrated, for example, in Refs. Shen and Piecuch, 2012; Lyakh et al., 2011, of other noniterative triples corrections to CCSD, including CCSD(T)Shen and Piecuch (2012); Lyakh et al. (2011) and CCSD(2)T,Shen and Piecuch (2012) to accurately describe the lowest $^{1}A_{g}(D_{2\text{h}})$ state of cyclobutadiene in the automerization barrier region is, in significant part, a consequence of the inability of all such methods to capture the coupling of the lower-rank $T_{1}$ and $T_{2}$ clusters with $T_{3}$, which in the vicinity of the TS geometry, where $T_{3}$ effects are large and nonperturbative, is big enough to substantially alter $T_{1}$ and $T_{2}$ amplitudes compared to their CCSD values. This means that to improve the quality of the $^{1}A_{g}(D_{2\text{h}})$ potential obtained with CR-CC(2,3), one should relax $T_{1}$ and $T_{2}$ clusters, adjusting them to the presence of $T_{3}$ correlations, prior to determining noniterative triples corrections. The CIPSI-driven CC($P$;$Q$) methodology allows us to do it in a computationally efficient manner, avoiding expensive CCSDT iterations, by incorporating the leading triply excited determinants identified with the help of the CIPSI runs using sufficiently large $N_{\text{det(in)}}>1$ values into the underlying $P$ spaces and correcting the resulting CC($P$) energies for the remaining $T_{3}$ effects using the $\delta(\mbox{$P$;$Q$})$ corrections.
@@
The situation with the lowest-energy $^{3}B_{1g}(D_{\rm 2h})$ potential is different. In this case, as shown in Table 2 and Figure 3, the $T_{3}$ correlation effects, estimated, after subtracting the CCSD energies from their CCSDT counterparts obtained with cc-pVDZ, at about (25)(24)(-25)-(-24) millihartree, barely vary with the automerization coordinate $\lambda$ and are accurately described by the CR-CC(2,3) approach, which reproduces the parent CCSDT energetics to within 60 microhartree across the entire $^{3}B_{1g}(D_{\rm 2h})$ PEC. Because of this very different behavior of CR-CC(2,3) compared to the lowest-energy singlet potential, which can be seen by comparing the $N_{\text{det(in)}}=1$ CC($P$;$Q$) PECs in the (b) and (d) panels of Figure 5, one ends up with a highly unbalanced description of the lowest singlet and triplet states of cyclobutadiene by the CR-CC(2,3) method, especially in vicinity of the barrier on the $^{1}A_{g}(D_{2\text{h}})$ potential. This is reflected in the errors characterizing the singlet–triplet gap values obtained with CR-CC(2,3), relative to CCSDT, which in the calculations using the cc-pVDZ basis set grow from $0.553$ kcal/mol at $\lambda=0$ to $9.222$ kcal/mol when the $\lambda=1$ square TS structure is considered [see the $N_{\text{det(in)}}=1$ CC($P$;$Q$) values of $\Delta E_{\text{S--T}}$ in Table 3; cf., also, Figures 4 and 5 (f)]. Once again, the main problem resides in the neglect of the coupling between $T_{1}$ and $T_{2}$ clusters and their higher-rank $T_{3}$ counterpart in the CR-CC(2,3) approach, which results in a poor description of the $^{1}A_{g}(D_{2\text{h}})$ potential in the vicinity of the TS geometry that propagates into the similarly poor $\Delta E_{\text{S--T}}$ gap values. To bring the results closer to those obtained with CCSDT, the input variable $N_{\text{det(in)}}$ that controls the CIPSI runs preceding the CC($P$) and CC($P$;$Q$) steps must be increased. The results of the CIPSI-driven CC(PP)/$c$-pVDZ and CC($P$;QQ)/$c$-pVDZ computations for the lowest $^{1}A_{g}(D_{2\text{h}})$ and $^{3}B_{1g}(D_{\rm 2h})$ potentials of cyclobutadiene and the gap between them using representative $N_{\text{det(in)}}>1$ values are discussed next.
@@
As shown in Table 1 and Figures 2 and 5 (b), the CC($P$;$Q$) computations for the lowest $^{1}A_{g}(D_{2\text{h}})$ state, using CIPSI Hamiltonian diagonalizations to identify the leading triply excited determinants for inclusion in the underlying $P$ spaces, display fast convergence toward CCSDT with $N_{\text{det(in)}}$, independent of the value of $\lambda$. In the case of the calculations performed using the cc-pVDZ basis set discussed here, with as little as 101,361101,361196,965196,965 $S_{z}=0$ determinants of the $A_{g}(D_{\rm 2h})$ symmetry in the terminal $|\Psi^{(\text{CIPSI})}\rangle$ wave functions generated by the inexpensive CIPSI runs using $N_{\text{det(in)}}=100,000$, which capture tiny fractions, on the order of 0.1–0.2%, of the $14,483,876$ $A_{g}(D_{\rm 2h})$-symmetric $S_{z}=0$ triples, the CC($P$;$Q$) method reduces the 0.848, 1.253, 2.021, 3.582, 7.008, and 14.636 millihartree errors relative to CCSDT obtained at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1 with CR-CC(2,3) to 0.431, 0.552, 0.776, 1.235, 0.628, and 3.539 millihartree, respectively. With the relatively small additional effort corresponding to $N_{\text{det(in)}}=250,000$, which results in 394,080394,080449,753449,753 $S_{z}=0$ determinants of the $A_{g}(D_{\rm 2h})$ symmetry in the final CIPSI diagonalization spaces and only 0.5–0.6% of all triples in the underlying $P$ spaces, the differences between the CC($P$;$Q$) and CCSDT energies of the lowest singlet state of cyclobutadiene at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1 decrease to 0.278, 0.272, 0.271, 0.300, 0.336, and 0.458 millihartree, respectively. Clearly, these are massive error reductions compared to the CR-CC(2,3) computations, especially in the barrier region, which highlight the effectiveness of our CIPSI-driven CC($P$;$Q$) strategy and the importance of relaxing $T_{1}$ and $T_{2}$ amplitudes in the presence of the leading $T_{3}$ contributions compared to their CCSD values prior to determining the noniterative corrections for the remaining $T_{3}$ effects. Similar comments apply to the improvements in the troublesome 13.788 millihartree NPE relative to CCSDT characterizing the $^{1}A_{g}(D_{2\text{h}})$ potential obtained in the CR-CC(2,3)/cc-pVDZ calculations offered by the CIPSI-driven CC($P$;$Q$) runs. When the CC($P$;QQ)/$c$-pVDZ approach using $N_{\text{det(in)}}=100,000$ is employed, the NPE relative to CCSDT characterizing the resulting $^{1}A_{g}(D_{2\text{h}})$ potential becomes 3.108 millihartree, which is a reduction of the corresponding CR-CC(2,3) NPE value by a factor of 4.4. The $N_{\text{det(in)}}=250,000$ CC($P$;$Q$) computations, which extract the lists of triples from the relatively inexpensive Hamiltonian diagonalizations in spaces that in the case of the cc-pVDZ basis set are 32–37 times smaller than the number of $T_{3}$ amplitudes used by CCSDT, reduce the 13.788 millihartree NPE characterizing the lowest-energy $^{1}A_{g}(D_{2\text{h}})$ potential obtained with CR-CC(2,3)/cc-pVDZ, relative to its CCSDT counterpart, to an impressively small value of 0.187 millihartree. This is a 74-fold reduction in NPE compared to CR-CC(2,3).